SOLVE THE EQUATION 4^(2x+1) = 2^(3x+6).
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

SOLVE THE EQUATION 4^2x+1=2^3x+6.
16x+1=8x+6
Subtract 8x from both sides to get:
8x=5
Divide both sides by 8 to get:
x=5/8
Cheers,
Stan H.

Answer by Edwin McCravy(20086)


     42x+1 = 23x+6

Rewrite 4 as (22)

  (22)2x+1 = 23x+6

Remove the parentheses on the left by
multiplying the inner exponent 2 by
the outer exponent 2x+1

  22(2x+1) = 23x+6

Now use the rule which says

if AB = AC with A > 0 and A ¹ 1,

    then B = C

So we have:

  2(2x + 1) = 3x + 6

Can you solve that? If not post again,
but I think you can.

Answer: x = 4

Edwin